OpenStudy (anonymous):

What is the smallest number that has exactly 12 factors?

OpenStudy (anonymous):

12!

OpenStudy (anonymous):

like factorial of 12, not 12

OpenStudy (anonymous):

Is it: \[2^{12}\]

OpenStudy (anonymous):

nop..i'll tell the ans later,,but thats not the ans..

OpenStudy (ragingsquirrel):

Isnt it 60?

OpenStudy (anonymous):

yep..60..

OpenStudy (ragingsquirrel):

Boosh.

OpenStudy (anonymous):

but i'd like to know the method..

OpenStudy (anonymous):

2.5.3.4.5.6.10.12.15.20.30.60 so this is what you meant :C, numbers of factors. not them multiplying each other or different factors ._.

OpenStudy (anonymous):

ofcorse thats what i meant,,question was quite clear,.anyways,,anyone can show some method to do it?

OpenStudy (anonymous):

by the way you repeated 5 and missed 1 !! :P

OpenStudy (anonymous):

different factors?

OpenStudy (anonymous):

yes..2.5.3.4.1.6.10.12.15.20.30.60 these are the 12 factors.. i want the method of knowing how 60 was smallest no. with exact 12 factors..

OpenStudy (anonymous):

tests and trials!

OpenStudy (anonymous):

well..i have 1 method,,i want to see if any one else posts any..

OpenStudy (anonymous):

so u mean 12 divisors not factors

OpenStudy (anonymous):

well isnt 4 a factor of 8 ?

OpenStudy (anonymous):

or is it the other way around? 8 a factor of 4 ?

OpenStudy (anonymous):

let me think

OpenStudy (anonymous):

are u familiar with prime factorization and divisors of a positive integer?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

12=3*2*2 three smallest factors of 12 are 2,2 and 3 3*2*2=(2+1)(1+1)(1+1) u just need to choose three different prime numbers with exponents 1,1,2 to make the number largest exponent for smallest prime number i choose 2 for 2 1 for 3 and 1 for 5 so n=2^2*3*5=60

OpenStudy (anonymous):

excellent !! :)